# CNN on ‘Top’: In Search of Scalable & Lightweight Image-based Jet Taggers

### v1

@@ 2.5 Training
Although we load the validation and test datasets into RAM as a whole, due to our choice of confining ourselves to a single mid-range PC with $64GB$ RAM, we use a type of data piping for the training dataset. We load $300,000$ training data at a time, which we will call a ‘round’ from now on. Each round is trained until the validation loss stops decreasing for 10 epochs. An image of the trained network is saved from each round and used in the next one. This is done for 15 (18) rounds for $32\times 32$ ($40\times 40$) image representation of jets to ensure that the entire dataset is adequately represented. For most of the models used in this work, we repeat the process, starting with the last trained network, and use the same rounds again. This way, the rounds mimic the behavior of epochs, while storing the datasets beforehand reduces computation time. We then evaluate the accuracy of each of these models on the test dataset, and the one with the highest accuracy is chosen as the best version of that specific model architecture. All further measurements are performed on this iteration of the model.
To ensure optimal computation time, we choose batch sizes depending on whether the memory dedicated to the graphics card can handle them without overflowing into the SWAP memory. For EffNet-S (64) a batch size of $2048$ is used. For all other networks, it is fixed at $4096$. We use the ADAM optimizer kingma2014adam with default settings ($\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-5}$, along with the default learning rate update schedule) to minimize the cross-entropy loss, where we wait for a fixed number of epochs (generally 10) for the last minimum loss to be updated; otherwise, we stop the training.
@@
We retrain the best network of each architecture nine additional times, with a representative subset of the dataset, to obtain the uncertainties arising from random weight initialization.
@@ 3 Results
Table (2) lists all single-number performance metrics for the networks explored in this analysis. The table is divided into three blocks: the first block contains all the information about the LeNet architecture, the second one about EffNet-S, and the third one lists some of the publicly available information about CNNs used previously as top-taggers for comparison. We have obtained results for LeNet-type networks with three different input resolutions: $28\times 28$, $32\times 32$, and $40\times 40$. Versions of EffNet-S are of two types: $40\times 40$ and $64\times 64$. Images are generated in two ways: $32\times 32$ images cropped to $28\times 28$ and $40\times 40$ images cropped to $32\times 32$, as explained in sec. 2.2. Each network is run on all images with cropped resolution (col. 3 of tab. 2) less than or equal to its input dimension (col. 2 of tab. 2). Networks with a specific architecture and input resolution optimized for maximum accuracy are used to create versions that can also process the global features.
@@
The smallest EffNet-S, while applied only to images, outperforms all LeNets, while being $1/8$-th the size of the largest LeNet and in terms of the total number of parameters. Its performance is best when the input images are $32\xrightarrow{Crop}28$, and it drops a little for the larger images. The possible reason is that the ‘MB-Conv’ blocks with bottlenecks in EffNet-S do not perform at the level of full convolutions when analyzing sparse pixel-level information, and networks with larger input arrays are needed for proper classification. Adding global features is beneficial here as well, but the effect is less prominent for larger networks. One possibility is that the extra MLP block after the last Aggregation layer contributes far less than the rest of the network. Analyzing higher input sizes and higher image resolutions can shed more light on this. We have shown the best metrics in each column in bold-face, for both the cases with and without global features.
Fig. (3) shows the signal (top) efficiency vs. the background (QCD) rejection curves for those representative models for which at least one of the metrics is highest among others (those in bold-face in table 2).
@@ 4 Conclusion
Jet tagging, a classification problem that aims to identify jets originating from different particles (top quarks, Higgs boson, gauge bosons, light quarks, or gluons), remains a challenging, albeit one of the most important tasks, especially in the context of a hadron collider like the LHC. Machine learning has now replaced traditional use of hand-crafted features to become an integral part of jet physics studies. Depending on the network used, jets are treated as images, sequences, trees or graphs. Existing literature shows that transformer based networks, which treat jets as particle clouds or graphs, perform the best. To achieve the highest possible accuracy, incorporating pairwise information is crucial. However, given the large number of pairwise interactions among the constituents of a jet, creating a fully connected graph becomes a computationally costly affair. CNNs, on the other hand, while being computationally low-cost, treat jets as images/arrays denoting the spatial arrangement of constituent particles in terms of energy, momentum, and mass; thus can utilize at least some local pairwise information. One can additionally incorporate global features of the jets to either directly to the images or to the network to enhance the jet identification accuracy. We have explored the possibility of combining these features and construct a computationally lightweight but efficient CNN-based network that can compete with other similar architectures.
@@
With traditional CNNs, larger/deeper networks are needed to extract maximal information from the jet images, as opposed to typically lightweight CNNs like LeNet. Scaled up CNNs, such as ResNets are computationally costly. Using recent advancements in the construction of CNNs, we have used lightweight versions of EfficientNet to circumvent this issue. Although these were built to work on higher resolution images, one can follow a set of prescribed scaling rules to construct a network applicable for low resolution jet images as well. We have performed a comparative study with and without externally provided global features. Results for LeNet are also provided for comparison.
@@
We observe that adding global features to the models, along with the images, results in better accuracy. Although this increase is somewhat monotonic with increasing image resolution for fully convolutional models, such as LeNet, the corresponding relationship for EffNet-S is not so straightforward. Increasing the input array size also improves the overall performance.
@@
This exploration leads to three observations: first, there needs to be a systematic architecture search for versions of scalable models such as EfficientNet, tuned for jet classification; second, adding global jet features, in addition to the constituent information, makes it easier for smaller models to perform better; last, but not the least, as jets can be represented in multiple types of data, one can combine the strengths of different network architectures and build a hybrid ensemble of lightweight networks capable of achieving high accuracy while working with various representations of the same jet, e.g., image, constituent four-momenta, and global jet features. This ‘mixture of experts’ approach has the potential for superior performance, leading to better signal-to-background discrimination. Exploring this idea is beyond the scope of this paper, but it is something that we plan to work on as a subsequent project.

### v2

@@ 2.5 Training
Although we load the validation and test datasets into RAM as a whole, due to our choice of confining ourselves to a single mid-range PC with $64GB$ RAM, we use a type of data piping for the training dataset. We load $300,000$ training data at a time, which we will call a ‘round’ from now on. Each round is trained until the validation loss stops decreasing for 10 epochs. An image of the trained network is saved from each round and used in the next one. This is done for 12 rounds for $35\times 35$ ($40\times 40$) image representation of jets to ensure that the entire dataset is adequately represented. For most of the models used in this work, we repeat the process, starting with the last trained network, and use the same rounds again. This way, the rounds mimic the behavior of epochs, while storing the datasets in advance reduces computation time. We then evaluate the accuracy of each model on the test dataset, and the model with the highest accuracy is saved as the best version of that specific architecture. All further measurements are performed on this iteration of the model.
To ensure optimal computation time, we choose batch sizes depending on whether the memory dedicated to the graphics card can handle them without overflowing into the SWAP memory. For EffNet-S versions 0, 11, and $2$, a batch size of $4096$ is used; for EffNet-S versions $3$ and $4$, these numbers are respectively $3072$ and $2560$. For all other networks, it is fixed at $4096$. We use the ADAM optimizer Kingma (2014) with default settings ($\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-5}$, along with the default learning rate update schedule) to minimize the cross-entropy loss, where we wait for a fixed number of epochs (generally 10) for the last minimum loss to be updated; otherwise, we stop the training.
@@
We retrain the best network of each architecture 9 additional times, with a representative subset of the dataset, to obtain the uncertainties arising from random weight initialization.
@@ 3 Results
Table (2) lists all single-number performance metrics for the networks explored in this analysis. The table is divided into three blocks: the first block contains all the information about the LeNet architecture, the second one about EffNet-S, and the third one lists some of the publicly available information about CNNs used previously as top-taggers for comparison. We have obtained results for LeNet-type networks with three two different input resolutions: $35\times 35$, and $40\times 40$. Versions of EffNet-S are of 5 types: 0 to $4$. Images are generated in two ways: $35\times 35$ images cropped to $28\times 28$ and $40\times 40$ images cropped to $32\times 32$, as explained in sec. 2.2. Each network is run on all images with cropped resolution (col. 3 of tab. 2) less than or equal to its input dimension (col. 2 of tab. 2). Networks with a specific architecture and input resolution optimized for maximum accuracy are used to create versions that can also process the global features.
@@
The smallest EffNet-S, while applied only to images, achieves accuracy similar to the largest LeNet, while being $1/8$-th in size (the total number of parameters). Its performance is best when the input images are $35\xrightarrow{Crop}28$, but it drops quite a lot for the larger images. Adding global features is beneficial here as well, but the effect is less prominent for larger networks. We have shown the best metrics in each column in bold-face, for both the cases with and without global features.
As was mentioned earlier in sec. 2.1, the inference timings were calculated in a separate machine with a weaker GPU. While networks consistently take more time for larger image sizes, all of the networks dealt with in this work have almost half the inference time of that of ResNeXt-50.
Fig. (3) summarizes the accuracies of table (2) in comparison to architecture complexity (number of net-parameters). Fig. (3(d)) shows the relative sizes of different versions of EffNet-S compared to the versions of LeNets used in this analysis as well as the larger and standard EfficientNets (B0 to B7), where the number of parameters is in log-scale. Even the largest EffNet-S is smaller than the smallest LeNet and all of these are an order of magnitude smaller than EffNet-Bs. In figures (3(a)) and (3(b)), accuracies of the nets are compared, with their model complexity in the $x$-axis. While fig. (3(a)) plots the best accuracies for each type of net, fig. (3(b)) shows the variation in it arising from random weight initialization. These employ 10 runs each, as mentioned at the end of the last section.
As far as accuracies are concerned and for only images as input, we see that the performance of LeNet does not change much (and remains within randomization error), with increasing complexity, for all image resolutions. The EffNet-S, on the other hand, performs worse for larger image sizes. For the larger variants, this loss in performance falls within the randomization error-bar, but overall, the performance not only worsens, but has larger variance for $40\xrightarrow{Crop}32$ images. The possible reason is that the ‘MB-Conv’ blocks with bottlenecks in EffNet-S do not perform at the level of full convolutions when analyzing sparse pixel-level information, and networks with larger input arrays are needed for proper classification.
While the addition of global features improves network performances by a margin, the accuracies become more or less independent of the network complexity across architectures, especially when randomization errors are considered. One possibility is that the extra MLP block after the last Aggregation layer contributes far less than the rest of the network, but the more probable explanation would be that the impact of global features overshadow the CNN-performance. Analyzing higher input sizes and higher image resolutions can shed more light on this. More work is needed regarding the integration of the global features in these networks.
Fig. (4) shows the signal (top) efficiency vs. the background (QCD) rejection curves for those representative models for which at least one of the metrics is highest among others (those in bold-face in table 2). This also shows how all EffNet-S nets outperform the best LeNet with just images as input (dashed curves), but with Global features put in, all of their performances are equivalent (solid, colored curves). We have also added the optimal possible ROC curve obtained from likelihood ratio analysis using generative models (ref. Geuskens:2024tfo; in thick black) to showcase the space allowed for future progress.
@@ 4 Conclusion
Jet tagging, a classification problem that aims to identify jets originating from different particles (top quarks, Higgs boson, gauge bosons, light quarks, or gluons), remains a challenging, albeit one of the most important tasks, especially in the context of a hadron collider like the LHC. Machine learning has now replaced traditional use of hand-crafted features to become an integral part of jet physics studies. Depending on the network used, jets are treated as images, sequences, trees or graphs. Existing literature shows that transformer-based networks, which treat jets as particle clouds or graphs, perform the best. To achieve the highest possible accuracy, incorporating pairwise information is crucial. However, given the large number of pairwise interactions among the constituents of a jet, creating a fully connected graph becomes a computationally costly affair. CNNs, on the other hand, while being computationally low-cost, treat jets as images/arrays denoting the spatial arrangement of constituent particles in terms of energy, momentum, and mass; thus can utilize at least some local pairwise information. One can additionally incorporate global features of the jets to either directly to the images or to the network to enhance the jet identification accuracy. We have explored the possibility of combining these features and construct a computationally lightweight but efficient CNN-based network that can compete with other similar architectures.
@@
With traditional CNNs, larger/deeper networks are needed to extract maximal information from the jet images, as opposed to typically lightweight CNNs like LeNet. Scaled-up CNNs, such as ResNets are computationally costly. Using recent advancements in the construction of CNNs, we have used lightweight versions of EfficientNet to circumvent this issue. Although these were built to work on higher resolution images, one can follow a set of prescribed scaling rules to construct a network applicable for low resolution jet images as well. We have performed a comparative study with and without externally provided global features. Results for LeNet are also provided as a benchmark for comparison.
@@
We observe that adding global features to the models, along with the images, results in better accuracy. Although this the increase in performance is somewhat monotonic with increasing image resolution for fully convolutional models, such as LeNet, the corresponding relationship for EffNet-S is not so straightforward; increasing the input array size also improves the overall performance. The performance of these models becomes somewhat independent of their architecture in the presence of the global features. When applied to jet-classification tasks, this begs for further inquiry into the interplay between global features and constituent-level information.
@@
This exploration leads to three observations: first, there needs to be a systematic architecture search for versions of scalable models such as EfficientNet, tuned for jet classification; second, adding global jet features, in addition to the constituent information, makes it easier for smaller models to perform better, but the exact way these features are fed-in needs further exploration; last but not least, as jets can have multiple types of representation, e.g., image, constituent four-momenta, and global jet features, one can combine the strengths of different network architectures and build a hybrid ‘mixture of experts’ ensemble of lightweight networks with the potential for superior performance, leading to better signal-to-background discrimination. Exploring these ideas is beyond the scope of this paper, but it is something that we plan to work on as a subsequent project.